A plane dives downward at an acceleration of 6.73 m/s2. If the plane’s initial speed was 350 km/h and its final speed was 400 km/h, how far did it travel in meters?
1
Expert's answer
2020-11-17T13:02:56-0500

Let the initial speed be "v_0 = 350km/h = \dfrac{350}{3.6}m/s", the final speed be "v_f = 400km/h = \dfrac{40}{3.6}m/s"., and the acceleration be "a = 6.73m\/s^2" . The time required for such increase in speed is:


"t = \dfrac{v_f - v_0}{a}"

During this time the plane travelled the following distance:


"d =v_0t + \dfrac{at^2}{2} = v_0\dfrac{v_f - v_0}{a} + \dfrac{(v_f - v_0)^2}{2a}\\
 \space \\
d = \dfrac{350}{3.6}\cdot \dfrac{\dfrac{400}{3.6}- \dfrac{350}{3.6}}{6.73} + \dfrac{(\dfrac{400}{3.6} - \dfrac{350}{3.6})^2}{2\cdot 6.73} \approx 214.97m"

Answer. 214.97m.